OpenStudy (anonymous):

an airplane is flying on a bearing of 335 degree at 530 mph. find the component form of the velocity of the airplane

OpenStudy (anonymous):

\[\sin25 = Vy/530, Vy = \sin25*530 = 224\]\[\cos25 = Vx/530, Vx = \cos25*530 = 480 m/s\]

OpenStudy (anonymous):

\[224 = 224 m/s\]

OpenStudy (anonymous):

\[\theta = 25, obtained from 360-335\]